# Classification of foliations via disintegration maps

### v1

@@ Abstract.
In this paper, we present a novel approach for analyzing the relationship between the supports of conditional measures and their geometric arrangement in Wasserstein space via the disintegration map. Our method establishes criteria to determine when such conditional measures arise from a metric measure foliation. Additionally, we provide a concrete example demonstrating how this framework can be applied to study deformations (or perturbations) of disintegration-induced foliations.
@@ 1. Introduction
The disintegration of measures offers a powerful framework for analyzing local and global probabilistic structures by decomposing a measure into conditional probabilities. Roughly speaking, given $X,Y$ measurable spaces, $\mu$ a measure on $X$, and a measurable map $\pi:X\to Y$, a disintegration allows us to break $\mu$ into a family of conditional measures $\{\mu_{y}\}_{y\in Y}$ on $X$, depending measurably on $y$, such that $\mu$ can be reconstructed by integrating $\mu_{y}$ against $\pi_{*}\mu$, the push-forward of $\mu$ by $\pi$. This decomposition is crucial in several areas, such as Probability Theory, Ergodic Theory, and Geometric Measure Theory, where it provides a rigorous way to handle conditional distributions.
@@
Disintegration theorems typically require topological assumptions to ensure existence and uniqueness of such decompositions, making them a powerful tool to define measures on fibered spaces or under projections, for instance. Although significant progress has been made in establishing the existence of conditional measures under several different hypothesis, most of the existing approaches neglect the geometric aspects of the underlying space.
@@
In a recent paper [PR25], the second and third authors used the intrinsic structures of probability spaces to tackle the disintegration of measures from a perspective which takes into accounting statistical and geometrical properties of dynamical systems. In particular, it is introduced the notion of disintegration maps, which are essentially functions that assign to each point $y\in Y$ its corresponding conditional measure $\mu_{y}$, providing a structured way to analyze a disintegration. We emphasize that these objects can trace geodesics in the space of probability measures, which correspond to the solutions of a time-evolving transport problem, and exhibit certain rigidity phenomena: if one of the endpoints measure in the geodesic path is absolutely continuous, then so are all. Therefore, it becomes evident that certain geometric properties of the base space control the regularity of the conditional measures.
@@
In this paper, we further explore disintegration maps by defining a notion of derivative for these objects. Our notion of derivative resembles a metric derivative, comparing in a appropriated way the distance between conditional measures and the distance between the supports of such measures. We remark that, in order to define this notion of derivative we do not require any differentiable structure on the base space. An interesting case arises in the context of Riemannian submersions, for example, since the distance between fibers equals the distance between base points. In this context, if each conditional measure has full support on its corresponding fiber, we show that the distance among conditional measures and the distance between base points are the same, and this isometry is captured by the derivative. We introduce a novel dense energy functional, similarly to what is done in geometric analysis. This notion is suited to quantify the relationship between these distances more accurately. We show that the minimization of this functional detect and characterize highly ordered geometric structures in the base space. More specifically, we show that the minimum of this energy functional corresponds to a metric measure foliation. That is, a foliation of the base space whose leaves are parallel to each other and where the distance between conditional measures on these leaves equals the distance between the leaves themselves. Moreover, this functional exhibits sensitivity to small deformations of the foliation, capturing fine geometric structures. We provide an explicit example in which we assess the sensitivity of the energy functional by inducing deformations in a metric measure foliation.
@@
The main contributions of this paper are presented in Section 4, where we introduce a novel notion of derivative for disintegration maps (Definition 4.1). We then construct an energy functional designed to quantify the relationship between the Wasserstein distance of conditional measures and the distance between their supports (Definition 4.2). Our main result is:
@@
The minimum of $\mathcal{E}(f)$, the energy of the disintegration map, is equals to one, if and only if, the supports of conditional measures form a metric measure foliation.
The precise statement shall be given in Theorem A.
The paper is organized as follows. Section 2 outlines the theoretical framework and necessary definitions. Section 3 reviews essential background on disintegration maps. Finally, Section 4 presents our core results, including a detailed example exploring the sensitivity of the energy functional to deformations of the leaves that support the conditional measures.
@@ 2. Theoretical framework
Let $X$, $Y$ be locally compact complete separable metric spaces. Consider $\pi:X\to Y$ a Borel map, and $\mu\in\mathcal{M}_{+}(X)$, where $\mathcal{M}_{+}(X)$ is the set of all finite positive Radon measures on $X$. Define $\nu:=\pi_{*}\mu$ in $\mathcal{M}_{+}(Y)$. Then, by [PR25, Theorem A] there exist measures $\mu_{y}\in\mathcal{M}_{+}(X)$, such that,
  1. (1)
    $y\mapsto\mu_{y}$ is a Borel map for $\nu$-almost every $y\in Y$;
@@
The family $\{\mu_{y}\}_{y\in Y}$ is called the disintegration of $\mu$ with respect to $\nu$.
in order to examine this family of probability measures, we endow $\mathscr{P}(X)$, the space of probability measures on $X$, with a metric structure. The framework of Wasserstein spaces provides a natural geometric setting for such investigations. The Wasserstein space of order $p$, for $p\in[1,\infty)$, is the set
$\mathscr{P}_{p}(X):=\Big{\{}\mu\in\mathscr{P}(X):\int d(x,\tilde{x})^{p}\mu(dx)<+\infty\Big{\}}$
$\tilde{x}\in X$ arbitrary, endowed with the Wasserstein distance
$W_{p}(\mu,\nu):=\Big{(}\inf\limits_{\gamma\in\Pi(\mu,\nu)}\int d(x_{1},x_{2})^{p}~d\gamma(x_{1},x_{2})\Big{)}^{\frac{1}{p}}.$
for $\mu$ and $\nu$ probability measures on $X$, where $\Pi(\mu,\nu)$ is the set of measures $\gamma\in\mathscr{P}(X\times X)$ with marginals $\mu$ and $\nu$.
Given a disintegration of $\mu$ with respect to $\nu$, we call disintegration map a function assigns to each $y\in Y$ a conditional measure $\mu_{y}\in\mathscr{P}_{p}(X)$.
@@
[PR25, Definition 4.1] Let $X$, $Y$ be locally compact complete separable metric spaces. Consider $\pi:X\to Y$ a Borel map, $\mu\in\mathcal{M}_{+}(X)$, and $\nu:=\pi_{*}\mu$. Given $\{\mu_{y}\}_{y\in Y}$ a disintegration of $\mu$ with respect to $\nu$, we define the disintegration map by
$f:D\subset Y$ $\to(\mathscr{P}_{p}(X),W_{p})$
@@
To give an example, consider $X=[0,1]\times[0,1]$ the unit square, $Y=[0,1]$, $\mu=\text{Leb}^{2}$ the two-dimensional Lebesgue measure on $X$, $\pi(x,y)=x$ the projection onto the first coordinate, and $\nu=\text{Leb}^{1}$ the one-dimensional Lebesgue measure on $Y$. The disintegration $\{\mu_{x}\}_{x\in[0,1]}$ of $\mu$ with respect to $\nu$ is given by:
$\mu_{x}=\text{Leb}^{1}\text{ on }\{x\}\times[0,1]\quad\text{(vertical line segment at }x\text{)}$
Explicitly, for any measurable $A\subseteq X$:
$\mu(A)=\int_{[0,1]}\mu_{x}(A)\,d\nu(x)=\int_{0}^{1}\text{Length}\big{(}A\cap(\{x\}\times[0,1])\big{)}\,dx$
where $\text{Length}(A\cap(\{x\}\times[0,1])):=\text{Leb}^{1}(y\in[0,1]~:~(x,y)\in A)$. In fact, note that the map $x\mapsto\mu_{x}(A)=\text{Length}(A\cap\{x\}\times[0,1])$ is measurable, for any $B\subseteq[0,1]$,
$\mu(A\cap\pi^{-1}(B))=\int_{B}\text{Length}\big{(}A\cap\{x\}\times[0,1]\big{)}\,dx$
and $\mu_{x}$ is supported on $\pi^{-1}(x)=\{x\}\times[0,1]$, satisfying the disintegration conditions.
Figure 1. Disintegration of $\text{Leb}^{2}$ with respect to $\text{Leb}^{1}$: the red line $\{x\}\times[0,1]$ carries $\mu_{x}=\text{Leb}^{1}$
@@
where $\mu_{x}$ is $\text{Leb}^{1}$ on $\{x\}\times[0,1]$. This is a basic example that fits into the broader framework of metric foliations, which is the primary focus of our study.
Let $(X,d)$ be a metric space, and $\mathcal{F}$ a partition of $X$ into closed subsets. We call $\mathcal{F}$ a foliation of $X$ and the elements of this partition are called leaves. If for every $F,F^{\prime}\in\mathcal{F}$ and every $x\in F$, $d(F,F^{\prime})=d(x,F^{\prime})$, where $d(F,F^{\prime})=\inf\{d(x,x^{\prime}):x\in F,x^{\prime}\in F^{\prime}\}$ and $d(x,F^{\prime})=d(\{x\},F^{\prime})$, then $\mathcal{F}$ is called a metric foliation. In case that each leaf is bounded, we say that $\mathcal{F}$ is bounded. Given a metric foliation $\mathcal{F}$ of $X$, define the equivalence relation:
@@
Consider the $X^{*}:=X/\sim$ the set of equivalence classes under (1) and the projection $\pi:X\to X^{*}$ onto $X^{*}$. We call $X^{*}$ the quotient space and $\pi$ the quotient map. Define a distance function $d^{*}$ on $X^{*}$ as
$d^{*}(y,y^{\prime}):=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))$ (2)
for $y,y^{\prime}\in X^{*}$. Note that $\pi$ is a submetry: $\pi(B(x,r))=B(\pi(x),r)$, where $B(x,r)$ is a ball centred at $x$ with radius $r$. In fact,
@@
Therefore, $\pi$ is 1-Lipschitz. In this notation, we define a metric measure foliation.
@@
Let $X$ be locally compact complete separable metric space, and $\mathcal{F}$ be a metric foliation of $X$. Consider $\mu\in\mathscr{M}_{+}(X)$, the quotient space $X^{*}$, and the quotient map $\pi:X\to X^{*}$. We call $\mathcal{F}$ a $p$-metric measure foliation, $p\in[1,\infty)$, if $\pi_{*}\mu$ is locally finite Borel measure on $X^{*}$, and
@@
for any $y,y^{\prime}\in Y$, where $\{\mu_{y}\}_{y\in Y}$ is a disintegration of $\mu$ with respect to $\pi_{*}\mu$.
@@
The general definition for a metric measure foliation requires $p=2$ and (3) only almost everywhere (see [GKMS18]). However, our definition is more restrictive, requiring (3) everywhere.
Note that, in this case, the disintegration map is an isometry. This is the most important characteristic of metric measure foliation for us. Although the definition holds for $p\in[1,\infty)$, our case of interest is $p=2$, since $\mathscr{P}_{2}(X)$ has a rich geometric structure [GRS25]. Moreover, if $\mathcal{F}$ is a 2-metric measure foliation, then
$W_{q}(\mu_{y},\mu_{y^{\prime}})=d^{*}(y,y^{\prime})=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))$
for any $q\in[1,\infty)$ and any $y,y^{\prime}\in X^{*}$, where $\{\mu_{y}\}_{y\in Y}$ is the disintegration of $\mu$ with respect to $\pi_{*}\mu$ [Kaz22, Proposition 3.11]. Henceforth, the term “metric measure foliation” will refer to what we have defined as a “2-metric measure foliation”.
A very important example of metric measure foliation is related to the action of isometry group [GR23]. Let $(X,d,\mu)$ be a metric measure space and $G$ a compact topological group. Let
@@
be an isometric action of $G$ on $X$. Suppose this action is metric measure isomorphic. That is, for every $g\in G$ the map $X\ni x\mapsto gx\in X$ is an isometry preserving the measure $\mu$. Consider $[x]$ the $G$-orbit of a point $x\in X$ and the quotient space $X/G$ endowed with the distance
@@
Consider $\pi:X\to X/G$ the projection map, that is, $\pi$ is given by $x\mapsto[x]$. The family $\mathcal{F}:=\{\pi^{-1}(y):y\in X/G\}$ is a metric measure foliation on $X$.
Other interesting examples arise from Riemannian submersions of weighted Riemannian manifolds [GKMS18], for instance.

### v2

# Characterization of foliations via disintegration maps

@@ Abstract.
In this paper, we present a novel approach for analyzing the relationship between the supports of conditional measures and their geometric arrangement in Wasserstein space via the disintegration map. Our method establishes criteria to determine when such conditional measures arise from a metric measure foliation. Additionally, we provide a example demonstrating how this framework can be applied to study perturbations of disintegration-induced foliations.
@@ 1. Introduction
The disintegration of measures offers a powerful framework for analyzing local and global probabilistic structures by decomposing a measure into conditional probabilities. Roughly speaking, given measurable spaces $X,Y$, a measure $\mu$ on $X$, and a measurable map $\pi:X\to Y$, a disintegration allows us to break $\mu$ into a family of conditional measures $\{\mu_{y}\}$ on $X$, depending measurably on $y$, such that $\mu$ can be reconstructed by integrating $\mu_{y}$ against $\pi_{*}\mu$, the push-forward of $\mu$ by $\pi$. This decomposition is crucial in several areas, such as Probability Theory, Ergodic Theory, and Geometric Measure Theory, where it provides a rigorous way to handle conditional distributions.
@@
Disintegration theorems typically require topological assumptions to ensure existence and uniqueness of such decomposition, making them a key instrument to define measures on fibered spaces or under projections, for instance. Although significant progress has been made in establishing the existence of conditional measures under several different hypothesis, most of the existing approaches neglect the geometric aspects of the underlying space.
@@
In a recent paper [PR25], the second and third authors used the intrinsic structures of probability spaces to tackle the disintegration of measures from a perspective which takes into account statistical and geometrical properties of dynamical systems. In particular, it introduced the notion of disintegration maps, which are essentially functions that assign to each point $y\in Y$ its corresponding conditional measure $\mu_{y}$, providing a structured way to analyze a disintegration. We emphasize that these objects can trace geodesics in the space of probability measures, which correspond to the solutions of a time-evolving transport problem, and exhibit certain rigidity phenomena: if one of the endpoints measure in the geodesic path is absolutely continuous, then so are all inner points of the geodesic. Therefore, it becomes evident that certain geometric properties of the base space control the regularity of the conditional measures.
@@
In this paper, we further explore disintegration maps by defining a notion of derivative for them. This derivative resembles a metric derivative, comparing the distance between conditional measures and the distance between the supports of such measures, without requiring a differentiable structure on the base space. Motivated by this derivative, we introduce a novel dense energy functional, similarly to what is done in Geometric Analysis, for cases in which the disintegration applies to every $y\in Y$. This functional characterizes highly ordered geometric structures in the base space: its minimum corresponds precisely to a metric measure foliation, that is, a foliation of the base space whose leaves are parallel to each other and where the distance between conditional measures on these leaves equals the distance between the leaves themselves.
Metric measure foliation arises in various contexts, including Riemannian submersions between weighted manifolds, isometry group actions, and warped products, for instance [GKMS18, Kaz22]. They are significant in Differential Geometry and Analysis due to their role in studying geometric properties, and providing a framework for convergence theory. One the one hand, metric foliations themselves represent a smooth partition of a Riemannian manifold into lower-dimensional submanifolds that are locally equidistant, similar to orbits of an isometric group action or parallel hypersurfaces [GG88]. On the other hand, metric measure foliations are instrumental in the convergence theory of metric measure spaces, for instance. Research indicates that properties like the curvature-dimension condition and the Cheeger energy functional can be preserved from the original space to its quotient via these foliations. This framework also aids in investigating the convergence of sequences of metric measure spaces with unbounded dimensions [Kaz22].
@@
The main contributions of this paper are presented in Section 4, where we introduce a notion of derivative for disintegration maps (Definition 4.1) and an energy functional $\mathcal{E}_{p}$ (Definition 4.2), designed to quantify the relationship between the Wasserstein distance of conditional measures and the distance between their supports. Our main result is:
@@
Given $f$ a disintegration map, the energy $\mathcal{E}_{p}(f)$ is equal to one precisely when the conditional measures have supports constituting a metric measure foliation.
The precise statement shall be given in Theorem A. The paper is organized as follows. In Section 2, we establish the theoretical framework and requisite definitions. Section 3 reviews the essential background on disintegration maps. Our core results are presented in Section 4. These include examples that motivate the hypothesis in Theorem A, and a further example applying the energy functional to analyze the evolution of a measure under a flow.
@@ 2. Theoretical framework
Let $X$, $Y$ be locally compact complete separable metric spaces. Consider a Borel map $\pi:X\to Y$, and $\mu\in\mathcal{M}_{+}(X)$, where $\mathcal{M}_{+}(X)$ is the set of all finite positive Radon measures on $X$. Define $\nu:=\pi_{*}\mu$ in $\mathcal{M}_{+}(Y)$. Then, by [PR25, Theorem A], there exist measures $\mu_{y}\in\mathcal{M}_{+}(X)$ such that
@@
The family $\{\mu_{y}\}$ is called the disintegration of $\mu$ with respect to $\nu$, and the measures $\mu_{y}$ are called conditional probabilities. In essence, this disintegration theorem provides a specific and constructive way to define conditional probabilities when the conditioning is based on a measurable function. This construction leads to more regular conditional probabilities, a property not universally guaranteed by the more abstract definition of conditional probability with respect to an arbitrary sub-$\sigma$-algebra.
In order to examine this family of probability measures, we endow the space $\mathscr{P}(X)$ of probability measures on $X$, with a metric structure. The framework of Wasserstein spaces provides a natural geometric setting for such investigations. The Wasserstein space of order $p$, for $p\in[1,\infty)$, is the set
$\mathscr{P}_{p}(X):=\Big\{\mu\in\mathscr{P}(X):\int d(x,\tilde{x})^{p}\mu(dx)<+\infty\Big\},$
with $\tilde{x}\in X$ arbitrary, endowed with the Wasserstein distance
$W_{p}(\mu,\nu):=\Big(\inf\limits_{\gamma\in\Pi(\mu,\nu)}\int d(x_{1},x_{2})^{p}~d\gamma(x_{1},x_{2})\Big)^{\frac{1}{p}},$
for probability measures $\mu$ and $\nu$ on $X$, where $\Pi(\mu,\nu)$ is the set of measures $\gamma\in\mathscr{P}(X\times X)$ with marginals $\mu$ and $\nu$. In this context, given a disintegration of $\mu$ with respect to $\nu$, by disintegration map we mean a function that assigns to each $y\in Y$ a conditional measure $\mu_{y}\in\mathscr{P}_{p}(X)$.
@@
[PR25, Definition 4.1] Let $X$, $Y$ be locally compact complete separable metric spaces. Consider a Borel map $\pi:X\to Y$, $\mu\in\mathcal{M}_{+}(X)$, and $\nu:=\pi_{*}\mu$. Given a disintegration $\{\mu_{y}\}$ of $\mu$ with respect to $\nu$, we define the disintegration map by
$f:Y$ $\to(\mathscr{P}_{p}(X),W_{p})$
@@
To give an example, consider the unit square $X=[0,1]\times[0,1]$, the two-dimensional Lebesgue measure $\mu=\text{Leb}^{2}$ on $X$, the projection onto the first coordinate $\pi(x,y)=x$, and the one-dimensional Lebesgue measure $\nu=\text{Leb}^{1}$ on $[0,1]$. A disintegration of $\mu$ with respect to $\nu$ is a family of measures $\mu_{x}=\text{Leb}^{1}\text{ on }\{x\}\times[0,1]$, with $x\in[0,1]$. Explicitly, for any measurable subset $A\subseteq X$,
$\mu(A)=\int_{[0,1]}\mu_{x}(A)\,d\nu(x)=\int_{0}^{1}\text{Length}\big(A\cap(\{x\}\times[0,1])\big)\,dx,$
where $\text{Length}(A\cap(\{x\}\times[0,1]))=\text{Leb}^{1}(\{y\in[0,1]:(x,y)\in A\})$. Moreover, the map $x\mapsto\mu_{x}(A)$ is measurable, and $\mu_{x}$ is supported on $\pi^{-1}(x)$, satisfying the disintegration conditions.
Figure 1. Disintegration of $\text{Leb}^{2}$ with respect to $\text{Leb}^{1}$: the red line $\{x\}\times[0,1]$ carries $\mu_{x}=\text{Leb}^{1}$.
@@
where $\mu_{x}$ is $\text{Leb}^{1}$ on $\{x\}\times[0,1]$. This is a basic example that fits into the broader framework of metric foliations, which is the primary focus of our study. Consider a metric space $(X,d)$, and a partition $\mathcal{F}$ of $X$ into closed subsets. We call $\mathcal{F}$ a foliation of $X$, and the elements of this partition are called leaves. If for every $\in\mathcal{F}, FF^{\prime}\in\mathcal{F}$, and every $x\in F$, holds $d(F,F^{\prime})=d(x,F^{\prime})$, where $d(F,F^{\prime})=\inf\{d(x,x^{\prime}):x\in F,x^{\prime}\in F^{\prime}\}$ and $d(x,F^{\prime})=d(\{x\},F^{\prime})$, then $\mathcal{F}$ is called a metric foliation. In case that each leaf is bounded, we say that $\mathcal{F}$ is bounded. Given a metric foliation $\mathcal{F}$ of $X$, define the equivalence relation:
@@
Consider the set of equivalence classes under (1) $X^{*}:=X/\sim$ and the projection $\pi:X\to X^{*}$ onto $X^{*}$. We call $X^{*}$ the quotient space and $\pi$ the quotient map. We define $d^{*}$ a distance on $X^{*}$ by
$d^{*}(y,y^{\prime}):=d(\pi^{-1}(y),\pi^{-1}(y^{\prime})).$ (2)
Note that $\pi$ is a submetry: $\pi(B(x,r))=B(\pi(x),r)$, where $B(x,r)$ is a ball centered at $x$ with radius $r$. Indeed,
@@
When the disintegration of $\mu$ with respect to a metric foliation $\mathcal{F}$ is such that the $p$-Wasserstein distance between two conditional measures $\mu_{y}$ and $\mu_{y^{\prime}}$ equals $d^{*}(y,y^{\prime})$, we say that $\mathcal{F}$ is a $p$-metric measure foliation. More precisely:
@@
Let $X$ be locally compact complete separable metric space, and $\mathcal{F}$ be a metric foliation of $X$. Consider $\mu\in\mathscr{M}_{+}(X)$, the quotient space $X^{*}$ as defined above, and the quotient map $\pi:X\to X^{*}$. We call $\mathcal{F}$ a $p$-metric measure foliation, for $p\in[1,\infty)$, if $\pi_{*}\mu$ is locally finite Borel measure on $X^{*}$, and
@@
for any $y,y^{\prime}\in X^{*}$, where $\{\mu_{y}\}$ is a disintegration of $\mu$ with respect to $\pi_{*}\mu$.
@@
The general definition for a metric measure foliation requires $p=2$ and (3) almost everywhere; see [GKMS18], for instance. However, our definition is more restrictive, requiring (3) everywhere. This restriction is incorporated into the definition in anticipation of the characterization established in Theorem A.
Although the definition holds for $p\in[1,\infty)$, our case of interest is $p=2$, since $\mathscr{P}_{2}(X)$ has a rich geometric structure [GRS25]. Moreover, if $\mathcal{F}$ is a 2-metric measure foliation, then $W_{q}(\mu_{y},\mu_{y^{\prime}})=d^{*}(y,y^{\prime})$, for any $q\in[1,\infty)$ and every $y,y^{\prime}\in X^{*}$, where $\{\mu_{y}\}$ is the disintegration of $\mu$ with respect to $\pi_{*}\mu$ [Kaz22, Proposition 3.11]. Henceforth, the term “metric measure foliation” will refer to what we have defined as a “2-metric measure foliation”.
A very important example of metric measure foliation is related to the action of isometry group [GR23]. Let $(X,d,\mu)$ be a metric measure space and consider a compact topological group $G$. Let
@@
be an isometric action of $G$ on $X$. Suppose this action is metric measure isomorphic. That is, for every $g\in G$ the map $X\ni x\mapsto gx\in X$ is an isometry preserving the measure $\mu$. Consider the $G$-orbit $[x]$ of a point $x\in X$ and the quotient space $X/G$ endowed with the distance
@@
Consider the projection map $\pi:X\to X/G$, that is, $\pi$ is given by $x\mapsto[x]$. The family $\mathcal{F}:=\{\pi^{-1}(y):y\in X/G\}$ is a metric measure foliation on $X$. Other interesting examples arise from Riemannian submersions of weighted Riemannian manifolds; see [GKMS18], for instance.